OpenStudy (anonymous):

Find the value of k if x^2 - 2x - k = 0 has one solution

Directrix (directrix):

If x^2 - 2x - k = 0 has one solution, then the discriminant must be zero.

Directrix (directrix):

The discriminant is b^2 - 4 ac.

Directrix (directrix):

In this problem, a = 1, b = -2, and c = -k.

Directrix (directrix):

Take the equation b^2 - 4 ac = 0 and replace b by (-2), a by (1), and c by (-k). You can then solve for k. Post what you get so we can check it together. @Chad123

Directrix (directrix):

What did you get?

OpenStudy (anonymous):

I got - 1 as my answer

OpenStudy (anonymous):

K= -1

OpenStudy (anonymous):

Thanks a lot

Directrix (directrix):

Yes. Let's see how it looks on the quadratic.

Directrix (directrix):

x^2 - 2x - k = 0 x^2 - 2x - (-1) = 0 x^2 - 2x + 1 = 0 (x -1)(x-1)=0 x = 1 or x = 1 ---> It works.